“SPECTRUM-ONLY” ASSIGNMENT OF CORE-PENETRATING AND CORE-NONPENETRATING RYD-BERG STATES OF CALCIUM MONOFLUORIDE

\underline{JEFFREY J. KAY}, DANIEL S. BYUN, JASON O. CLEVENGER^{a}, XING JIANG, VLADIMIR S. PETROVIĆ, ROBERT SEILER^{b}, JONATHAN R. BARCHI, Department of Chemistry, Massachusetts Institute of Technology, Cambridge, MA 02139; ANTHONY J. MERER, Department of Chemistry, University of British Columbia, Vancouver, British Columbia, Canada V6T 1Z1; and ROBERT W. FIELD, Department of Chemistry, Massachusetts Insititue of Technology, Cambridge, MA 02139.

Rydberg states of calcium monofluoride in the n*=17–20 region have been observed by ionization-detected optical-optical double resonance spectroscopy via the D^{2}Σ v=1 intermediate state. We describe “spectrum-only” assignment procedures that can be used to assign optical-optical double resonance spectra of core-penetrating and core-nonpenetrating Rydberg states using only information contained in the spectrum rather than predictions derived from an effective Hamiltonian model. Using these methods we have assigned all members of the six core-penetrating Rydberg series and several components of f and g core-nonpenetrating Rydberg states in the n*=17–20 region, and have observed several perturbations between core-penetrating and core-nonpenetrating states. The analysis of these perturbations may enable disentanglement of the dipole, quadrupole, and higher-order ion-core multipole moments.

aPresent Address: Applied Materials, 3050 Bowers Avenue, Santa Clara, CA 95054

bPresent Address: Physical Chemistry Lab, ETH, CH-8093 Zurich, Switzerland